Find the variance of t? t 0 1 2 3 4 f(t) 1/9 2/9 3/9 2/9 1/9
1
Expert's answer
2021-09-06T16:32:58-0400
"\begin{matrix}
 t & & 0 & 1 & 2 & 3 & 4 \\
 f(t) & & 1/9 & 2/9 & 3/9 & 2/9 & 1/9
\end{matrix}"

"E(T)=\mu=\dfrac{1}{9}(0)+\dfrac{2}{9}(1)+\dfrac{3}{9}(2)+\dfrac{2}{9}(3)+\dfrac{1}{9}(4)"

"=2"

"E(T^2)=\dfrac{1}{9}(0)^2+\dfrac{2}{9}(1)^2+\dfrac{3}{9}(2)^2+\dfrac{2}{9}(3)^2+\dfrac{1}{9}(4)^2"

"=\dfrac{16}{3}"

"Var(T)=E(T^2)-(E(T))^2=\dfrac{16}{3}-(2)^2=\dfrac{4}{3}"

"Var(T)=E((T-\mu)^2)"

"=\dfrac{1}{9}(0-2)^2+\dfrac{2}{9}(1-2)^2+\dfrac{3}{9}(2-2)^2+\dfrac{2}{9}(3-2)^2"

"+\dfrac{1}{9}(4-2)^2=\dfrac{4}{3}"



"Var(T)=\dfrac{4}{3}"